OpenStudy (anonymous):

Can anyone demonstrate how to add two fractions with similar denominators and explain how to add them? And then the further steps to add two fractions with dissimilar denominators and the process demonstrated with an example?

OpenStudy (anonymous):

also known as rational expressions

OpenStudy (anonymous):

1/ similar denominator: just add / subtract the numerator, keep the denominator as it is ( just it is, not add/ subtract them) for example \[\dfrac{2}{5}+\dfrac{1}{5}=\dfrac{2+1}{5}=\dfrac{3}{5}\] got me so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the same with subtract, for example: \[\dfrac{2}{5}-\dfrac{1}{5}=\dfrac{2-1}{5}=\dfrac{1}{5}\] ok?

OpenStudy (anonymous):

Now dissimilar denominators: You have to make them have the same denominator. The easiest way is cross timing the denominator. Let me illustrate: \[\dfrac{2}{5}+\dfrac{1}{4}=\dfrac{2}{5}*\color{red}{\dfrac{4}{4}}+\dfrac{1}{4}*\color{blue}{\dfrac{5}{5}}=\dfrac{8}{20}+\dfrac{5}{20}\] Now, you have the form of add/subtract 2 fractions with the same denominator, right??

OpenStudy (anonymous):

yes so the final answer is 13/20 right?

OpenStudy (anonymous):

yup